Condensed balance sheet and income statement data for Flounder Corporation are presented here Flounder Corporation Balance Sheets December 31 2020 2019 2018 Cash Accounts receivable (net) Other current assets Investments Plant and equipment (net) $ 33,000 $23,000 $21,000 53,000 92,000 57,000 500,000 48,000 97,000 72,000 370,000 51,000 66,000 47,000 358,000 $735,000 $610,000 $543,000 Current liabilities Long-term debt Common stock, $10 par Retained earnings $ 88,000 83,000 $73,000 87,000 312,000 128,000 $735,000 $610,000 $543,000 147,000 322,000 178,000 52,000 302,000 116,000Flounder Corporation Income Statements For the Years Ended December 31 2020 2019 Sales revenue Less: Sales returns and allowances Net sales Cost of goods sold Gross profit Operating expenses (including income taxes) Net income $742,000 $602,000 33,000 569,000 352,000 217,000 153,000 $.89,000 $ 64,000 43,000 699,000 427,000 272,000 183,000 Additional information 1. The market price of Flounders common stock was $7.00, $7.50, and $8.50 for 2018, 2019, and 2020, respectively 2. You must compute dividends paid. All dividends were paid in cash Compute the following ratios for 2019 and 2020, (Round answers to 2 decimal places, eg. 1.83 or 1.83%.)2020 2019 Profit margin 12.73 % 11.24 % ross profit rate 38.91 % 38.131 % Asset turnover 1.03 times 0.98 times Earnings per share Price-earnings ratio Payout ratio 2.76 2.05 3.07 times 3.65 times 81.4111 % 27.861 % 43.7711 % Debt to assets ratio 31.97 %
Answer #1

Solution :

Earning per share = Net Income /Weighted Average outstanding shares

Weighted average outstanding shares:

2020 = (32200 + 31200) / 2 = 31700 shares

2019 = (31200 + 30200)/2 = 30700 shares

Earning per share:

2020 = $89,000/31700 = $2.81 per share

2019 = $64,000/30700 = $2.08 per share

Dividend paid = Beginning retained earnings + Net Income - Ending retained earnings

2019 = $116,000 + $64,000 - $128,000 = $52,000

2020 = $128,000 + $89,000 - $178,000 = $39,000

Payout ratio = Dividend paid / Net Income

2020 = $39,000 / $89,000 = 43.82%

2019 = $52,000 / $64,000 = 81.25%